the perimeter of a rectangular pool is 130 yards. 3 times the length is equal to 10 times the width. what is the length, width, and area of the pool?
Answer by mananth(12270)

let width be x
10W=3L
L=10W/3
...
length = 10x/3
perimeter = 130 yards
..
Perimeter = 2*(L+W)
2*(10x/3+x)=130
2*(13x/3)=130
13x/3 =65
13x= 195
x=15 yards the width
length = 10x/3
=150/3
= 50 yards the length